40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9509     Entry time: Sat Dec 21 01:54:04 2013
Author: Koji 
Type: Configuration 
Category: LSC 
Subject: LSC Whitening for the DCPDs/CM servo replaced 

The previous LSC whitening filters for the DCPDs were in an unknown state (although the transfer functions were actually measured and fit a while ago)
They had no DC gain control and some of the channels had modifications.

To make the setup clear, the filter module was replaced with the spare module without any modification.

The channels are now respoding to the whitening gain switches. So far there is no screen for the new  whitening gains yet.

Also I found that POX11 DC cable has not been connected. Now it is connected.

ELOG V3.1.3-